OpenStudy (anonymous):

What property is shown by this equation: (8 + 3) + 4 = 8 + (3 + 4 A. 1.83 x 10¹² B. 1.83 x 10⁻¹² C. 183 x 10¹° D. 1.83 x 10¹° @heavenfun13 @King_Trae_OKC

OpenStudy (anonymous):

(8+3) +4 = 8 + (3+4) 11+4 = 8 + 7 15=15

OpenStudy (anonymous):

Thats all ik... i dont see how its gonna be those answers tho

OpenStudy (anonymous):

Okay thanks anyways